# Unbiased Evaluation of Large Language Models from a Causal Perspective

### v1

@@ Abstract
Hikvision Research Institute
footnotetext: Equal contribution. Correspondence to: Meilin Chen (merlinarer@gmail.com), Liang Ma (maliang6@hikvision.com).
@@ 1 Introduction
Figure 2: Model bias visualizations. Left: $\mathcal{R}_{OC}$ vs. Strength on ARC-C and MMLU datasets. Right: $\mathcal{R}_{UC}$ vs. Strength on ARC-C and MMLU datasets. The point where strength equal zero represents the original datasets. Mistral, Yi, Llama, Qwen represents Mistral-Large-2411, Yi1.5-34B-Chat, Llama3.1-70B-Instruct and Qwen2.5-72B-Instruct, respectively. For Agents-as-an-Evaluator, we observe a significant increase in $\mathcal{R}_{OC}$ with growing strength, while $\mathcal{R}_{UC}$ remains relatively stable, indicating the existence of model bias. Compared with Agents-as-an-Evaluator, our Unbiased Evaluator remains relatively stable on both $\mathcal{R}_{OC}$ and $\mathcal{R}_{UC}$.
We termed this paradigm as Agents-as-an-Evaluator. Different from previous LLM-as-a-Judge (Zheng et al., 2023), which operates on the evaluation side by solely determining whether something falls within the scope of a given rule, Agents-as-an-Evaluator extends beyond assessment to the generation side, where LLMs actively contribute to the design of the very questions involved in the task. Considering that prior works have revealed that LLM-as-a-Judge posses certain biases(Blodgett et al., 2020; Ahn et al., 2022; Ferrara, 2023; Gallegos et al., 2024), compared to LLM-as-a-Judge, the critical step of question generation in Agents-as-an-Evaluator may introduce a greater potential for bias. Therefore, an important question rises:
@@ 3.1 Theoretical Analysis
$\vspace{-2em}\epsilon(\hat{\phi})=\mathbb{E}[\hat{\phi}]-\phi\vspace{-1em}$ (1)
@@
Proposition 3.1. (Proof in Appendix C.1) Given a new designed evaluation protocol, which transforms original benchmark $D$ into $D^{\prime}$. Then, the strength of bias with the new evaluation protocol can be decomposed into three terms: original, related and independent term.
@@
$+\underbrace{2\mathbb{E}[\epsilon(\hat{\phi}_{D})]\mathbb{E}[\Delta]+\mathbb{E}[\Delta^{2}]}_{\texttt{independent}}\vspace{-3em}$ (2)
@@ 3.3 Probing Task
where $\mathcal{S}_{ijk}$ is the confidence score assessed by $\theta_{k}$ for question $x^{\prime}_{ij}$. The score $\mathcal{S}_{ijk}$ lies within the range $[1,10]$, where a higher score indicates that the model has greater confidence in the ground truth label being the correct answer, while a lower score suggests stronger confidence that the ground truth label is not the correct answer. Detailed prompt is presented in Appendix  A.
@@
Table 2: Bags of Atomic Interventions (BOAT). This is a demo case for a clear demonstration. For further details, please refer to Sec. 5.1.
$\mathcal{T}$ positions interventions examples
original intervened label
Is 9.8 bigger than 9.11?
A: True B: False
question Distractor Question Is 9.8 bigger than 9.11? 1+1=? A
answer Answer Removal A: 2 B: False N
Option Shuffling A: False B: True B
Label Replacement I: True II: False I
Math overall Question Jitter Answer this question.
1+1=?
Judge True (T) or False (F) for given possible answer. 1+2=2? F
@@ 4.1 General formulation
Traditional evaluation methods, as depicted in Figure 3 (a), rely on static and fixed inputs, i.e., input variables, to investigate causal effects. However, this rigid framework may inadvertently introduce contamination issues, limiting the validity and robustness of the causal analysis. In contrast, our proposed approach, termed the Unbiased Evaluator, adopts a more dynamic and adaptive strategy, as shown in Figure 3 (b). Specifically, given input variables and ground truth label pairs $\{\{A,B,C,D\},R\}$ for a task $\mathcal{T}$, we enhance the evaluation by augmenting variables through carefully designed Bags Of Atomic InTerventions (BOAT), i.e. new input variable configurations, such as $\{\{\hat{A},B,\hat{C},D\},\hat{R}\}$ and $\{\{\hat{B},\hat{C},E\},\hat{R}\}$, are intervened with $\mathcal{F}(\cdot|\mathcal{T})$. More intervened examples are presented in Appendix D.
Our proposed approach not only mitigates potential contamination by incorporating diverse interventions but also offers an interpretable framework for evaluating LLMs. By examining counterfactual scenarios, such as comparing the effects of $\hat{A}$ versus $A$, our method fosters a deeper and more transparent understanding of model performance. We demonstrate this in the following experiments procedures.
@@
open-source Qwen2.5-72B-Instruct 94.33 85.69 $\pm$ 0.86 8.64 84.07 69.56 $\pm$ 0.39 14.51 98.41 88.86 $\pm$ 0.89 9.55
@@ 4.2 Bags of Atomic Interventions
Based on the general formulation described above, we provide the details of our designed Bags of Atomic Interventions in Table 2. Specifically, we focus on two widely used tasks: Multiple Choice Questions (MCQ) and Mathematics. For Multiple Choice Questions, we have designed six atomic interventions targeting different intervention positions within the task.
@@ 5.1 Experimental Setup
Implementation Details. We build our method on the widely-used OpenCompass (Contributors, 2023) evaluation framework. Following the approach in (Zhu et al., 2024), we set the generation temperature to 0 for all models and cap the output length at a maximum of 1000 tokens. All evaluations are conducted in a 5-shot setting, with results averaged over 5 runs. Moreover, the interventions in BOAT are not randomly combined but follow specific constraints. We regulate the probability of each intervention to ensure balance. When applying a binary transformation to questions, modifications involving phrases such as “which” or “following” were excluded. Furthermore, during the Answer Removal process, we ensured that the answers extracted from different questions were not identical. For additional details, please refer to Appendix B.
@@ 5.7 Human Verification
To ensure the correctness of proposed method, we randomly selected 300 samples from the MMLU and 100 samples from the ARC-C and GSM8K, 500 questions in total. 9 human experts (with bachelor or higher degree) are divided into 3 groups, each with 3 person. They were asked to judge the following question: Whether the answers to the intervened questions are correct. As shown in Table E, the human verification demonstrates an overall accuracy rate of 99.3%, 99.9% and 99.7% for ARC-C, MMLU and GSM8K, respectively, indicating the effectiveness of our methodology.
@@ 7 Impact Statement
Appendix A Prompt in Probing Task
@@
Appendix B Implementation Details
B.1 Hyperparameter
@@
B.2 BOAT constraints
As shown in 4, there are constraints among different interventions, and not all interventions can be combined arbitrarily. For example, the Answer Removal can only be introduced when the Distractor Hint intervention exists. Moreover, the binary transformation intervention cannot be combined with other interventions.
Please note that the instructions will change depending on the interventions. The detailed instructions are presented in 5.
Table 4: The BOAT constraints are represented in a matrix where each cell indicates whether the second intervention (column) can be applied when the first intervention (row) is introduced. A check mark (✓) denotes that the second intervention can be combined with the first one, while a black triangle ($\blacktriangle$) indicates that the second intervention is a required addition.
First/Second Intervention Distractor Hint Distractor Question Answer Removal Option Shuffling Label Replacement Binary Transformation
Distractor Hint $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Distractor Question $\checkmark$ $\checkmark$ $\checkmark$
Answer Removal $\blacktriangle$ $\checkmark$ $\checkmark$
Option Shuffling $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Label Replacement $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Binary Transformation
Table 5: The detailed instructions under different interventions.
@@ B.2 BOAT constraints
Distractor Hint If there is no correct answer in the options, please reply with N.
Distractor Question Here are two questions and only one of them corresponds to the options. Please select the correct answer.
Answer Removal Here is a multiple choice question.
Option Shuffling Here is a multiple choice question.
Label Replacement Here is a multiple choice question.
Binary Transformation The following are true/false questions. If the answer is correct, please reply with T, otherwise reply with F.
Appendix C Proof of Proposition
C.1 Proof of Proposition 3.1
@@ Appendix C Proof of Proposition
Appendix D Intervened Examples
@@ F T F F
Appendix E Human Verification
ARC-C MMLU GSM8K Average
Group 1 1.000 0.997 0.990 0.996
Group 2 0.990 1.000 1.000 0.997
Group 3 0.990 1.000 1.000 0.997
Average 0.993 0.999 0.997 /

### v2

@@ 1 Introduction
Figure 2: Model bias visualizations. Left: $\mathcal{R}_{OC}$ vs. Strength on ARC-C and MMLU datasets. Right: $\mathcal{R}_{UC}$ vs. Strength on ARC-C and MMLU datasets. Strength refers to the probability defined in Equation 3.3. A higher strength value indicates a greater proportion of ”processed” samples within the dataset (”process” denotes rephrasing and BOAT in Agents-as-an-Evaluator and Unbiased Evaluator, respectively). The point where strength=0 represents the original datasets. Mistral, Yi, Llama, Qwen represents Mistral-Large-2411, Yi1.5-34B-Chat, Llama3.1-70B-Instruct and Qwen2.5-72B-Instruct, respectively. For Agents-as-an-Evaluator, we observe a significant increase in $\mathcal{R}_{OC}$ with growing strength, while $\mathcal{R}_{UC}$ remains relatively stable, indicating the existence of model bias. Compared with Agents-as-an-Evaluator, our Unbiased Evaluator remains relatively stable on both $\mathcal{R}_{OC}$ and $\mathcal{R}_{UC}$.
We termed this paradigm as Agents-as-an-Evaluator. Formally, Agents-as-an-Evaluator refers to an LLM-based evaluation paradigm in which LLMs (or Agents) not only assess responses but also actively contribute to generating evaluation criteria and questions. Different from previous LLM-as-a-Judge (Zheng et al., 2023), which operates on the evaluation side by solely determining whether something falls within the scope of a given rule, Agents-as-an-Evaluator extends beyond assessment to the generation side, where LLMs actively contribute to the design of the very questions involved in the task. Considering that prior works have revealed that LLM-as-a-Judge posses certain biases(Blodgett et al., 2020; Ahn et al., 2022; Ferrara, 2023; Gallegos et al., 2024), compared to LLM-as-a-Judge, the critical step of question generation in Agents-as-an-Evaluator may introduce a greater potential for bias. Therefore, an important question rises:
@@ 3.1 Theoretical Analysis
$\epsilon(\hat{\phi})=\mathbb{E}[\hat{\phi}]-\phi$ (1)
@@
Proposition 3.1. (Proof in Appendix D.1) Given a new designed evaluation protocol, which transforms original benchmark $D$ into $D^{\prime}$. Then, the strength of bias with the new evaluation protocol can be decomposed into three terms: original, related and independent term.
@@
$+\underbrace{2\mathbb{E}[\epsilon(\hat{\phi}_{D})]\mathbb{E}[\Delta]+\mathbb{E}[\Delta^{2}]}_{\texttt{independent}}$ (2)
@@ 3.3 Probing Task
where $\mathcal{S}_{ijk}$ is the confidence score assessed by $\theta_{k}$ for question $x^{\prime}_{ij}$. The score $\mathcal{S}_{ijk}$ lies within the range $[1,10]$, where a higher score indicates that the model has greater confidence in the ground truth label being the correct answer, while a lower score suggests stronger confidence that the ground truth label is not the correct answer. Detailed prompt is presented in Appendix  B.
@@ 4.1 General formulation
Traditional evaluation methods, as depicted in Figure 3 (a), rely on static and fixed inputs, i.e., input variables, to investigate causal effects. However, this rigid framework may inadvertently introduce contamination issues, limiting the validity and robustness of the causal analysis. In contrast, our proposed approach, termed the Unbiased Evaluator, adopts a more dynamic and adaptive strategy, as shown in Figure 3 (b). Specifically, given input variables and ground truth label pairs $\{\{A,B,C,D\},R\}$ for a task $\mathcal{T}$, we enhance the evaluation by augmenting variables through carefully designed Bags Of Atomic InTerventions (BOAT), i.e. new input variable configurations, such as $\{\{\hat{A},B,\hat{C},D\},\hat{R}\}$ and $\{\{\hat{B},\hat{C},E\},\hat{R}\}$, are intervened with $\mathcal{F}(\cdot|\mathcal{T})$. More intervened examples are presented in Appendix E.
Our Unbiased Evaluator aims to assess whether models can genuinely answer a question correctly by employing causal interventions that align with human recognition. Unbiased Evaluator not only mitigates potential contamination by incorporating diverse interventions but also offers an interpretable framework for evaluating LLMs. By examining counterfactual scenarios, such as comparing the effects of $\hat{A}$ versus $A$, our method fosters a deeper and more transparent understanding of model performance. We demonstrate this in the following experiments procedures.
@@
Qwen2.5-72B-Instruct 94.33 85.69 $\pm$ 0.86 8.64 84.07 69.56 $\pm$ 0.39 14.51 98.41 88.86 $\pm$ 0.89 9.55
@@ 4.2 Bags of Atomic Interventions
Based on the general formulation described above, we provide a demo in Table 6 for a clear demonstration of our designed Bags of Atomic Interventions. Specifically, we focus on two widely used tasks: Multiple Choice Questions (MCQ) and Mathematics. For Multiple Choice Questions, we have designed six atomic interventions targeting different intervention positions within the task.
@@ 5.1 Experimental Setup
Implementation Details. We build our method on the widely-used OpenCompass (Contributors, 2023) evaluation framework. Following the approach in (Zhu et al., 2024), we set the generation temperature to 0 for all models and cap the output length at a maximum of 1000 tokens. All evaluations are conducted in a 5-shot setting, with results averaged over 5 runs. Moreover, the interventions in BOAT are not randomly combined but follow specific constraints. We regulate the probability of each intervention to ensure balance. When applying a binary transformation to questions, modifications involving phrases such as “which” or “following” were excluded. Furthermore, during the Answer Removal process, we ensured that the answers extracted from different questions were not identical. For additional details, please refer to Appendix C.
@@ 6.1 Bias Analysis
6.2 Contamination Analysis
Our Unbiased Evaluator aims to assess whether models can genuinely answer a question correctly by employing causal interventions that align with human recognition. Therefore, Unbiased Evaluator provides a more accurate measure of a model’s true and robust performance on a given benchmark by eliminating performance inflation caused by data contamination. The decline of accuracy in Table 2 actually reflects the decrease of contamination. To validate this, following (Yang et al., 2023), we further provide an additional fine-tuning ablation study. Specifically, we fine-tune Llama2-13B on the original samples from the MMLU test set and evaluate it on MMLU test set under two conditions: with and without our Unbiased Evaluator.
Table 3: Contamination analysis. We fine-tune Llama2-13B on the original samples from the MMLU test set and evaluate it on MMLU test set under two conditions: with and without our Unbiased Evaluator.
train set w/o w/
Llama2-13B 55.6 33.7
Llama2-13B + original test set 96.6 37.1
Results in 3 highlight that our Unbiased Evaluator provides a more rigorous assessment of benchmark contamination. Even when trained directly on the original test set, the model struggles to perform well under the Unbiased Evaluator, suggesting that it effectively mitigates data contamination and ensures a more robust evaluation.
6.3 Evaluation Reliability
Our Unbiased Evaluator demonstrates a significantly stronger alignment with human expert judgments. Since obtaining comprehensive expert evaluations across multiple models is both costly and impractical, we instead benchmark our method against LiveBench (White et al., 2025), a continuously updated and widely recognized evaluation platform. Specifically, we compute both Pearson and Kendall correlation coefficients between our averaged results of Table 2 and the global average scores reported in the latest LiveBench release (2024-11-25). To ensure a fair comparison, we exclude two models—GPT-4-Turbo and Yi1.5-34B-Chat—as they are not covered in the corresponding LiveBench update.
Table 4: Evaluation Reliability. We compute both Pearson and Kendall correlation coefficients between our averaged results of Table 2 and the global average scores reported in the latest LiveBench release.
Pearson Kendall
Vanilla 0.918 0.600
Unbiased Evaluator 0.949 1.000
Results in Table 4 confirm that our method aligns more closely with LiveBench. Notably, it achieves a perfect ranking correlation with LiveBench (as measured by Kendall), a significant improvement over baseline. Unlike LiveBench, which covers diverse tasks and requires substantial resources to update questions regularly, ours leverages existing benchmarks and requires almost no additional resources.
@@ 6.7 Human Verification
To ensure the correctness of proposed method, we randomly selected 300 samples from the MMLU and 100 samples from the ARC-C and GSM8K, 500 questions in total. 9 human experts (with bachelor or higher degree) are divided into 3 groups, each with 3 person. They were asked to judge the following question: Whether the answers to the intervened questions are correct. As shown in Table 6.7, the human verification demonstrates an overall accuracy rate of 99.3%, 99.9% and 99.7% for ARC-C, MMLU and GSM8K, respectively, indicating the effectiveness of our methodology.
ARC-C MMLU GSM8K Average
Group 1 1.000 0.997 0.990 0.996
Group 2 0.990 1.000 1.000 0.997
Group 3 0.990 1.000 1.000 0.997
Average 0.993 0.999 0.997 /
7 Future Work
Our proposed method can be easily extended towards open-ended tasks. Most tasks (e.g., multiple-choice, math), as demonstrated in this paper, inherently follow natural rules in either the questions or answers, and rule-based interventions can be automatically applied. The other small percentage of tasks, can use a debiased Agents-as-Evaluator version. Concretely, our study has revealed the data and model biases of previous version, inspiring two designs to mitigate them: (1) Cross-generation: to reduce model bias, we can break down question generation into multiple chunks, using different models for each. (2) Cross-checking: multiple advanced models can be used to cross-check the output to mitigate data bias and enhance quality. Overall, our method is easily scaled to most tasks, and our insights will provide valuable inspiration for future advancements in evaluation methodologies. We leave these directions to future work.
@@ 8 Conclusions
Appendix A Demo Case of BOAT
Table 6: Bags of Atomic Interventions (BOAT). This is a demo case for a clear demonstration. For further details, please refer to Sec. 5.1. For each intervention, we present a simple question to show its effect. The ”intervened” column only displays the part that has changed, highlighted in the same color as the original part it replaces. The other parts of the content remain unchanged.
$\mathcal{T}$ positions interventions examples
original intervened label
Is 9.8 bigger than 9.11?
A: True B: False
question Distractor Question Is 9.8 bigger than 9.11? 1+1=? A
answer Answer Removal A: 2 B: False N
Option Shuffling A: False B: True B
Label Replacement I: True II: False I
Math overall Question Jitter Answer this question.
1+1=?
Judge True (T) or False (F) for given possible answer. 1+2=2? F
Appendix B Prompt in Probing Task
@@
Appendix C Implementation Details
C.1 Hyperparameter
@@
C.2 BOAT constraints
As shown in 7, there are constraints among different interventions, and not all interventions can be combined arbitrarily. For example, the Answer Removal can only be introduced when the Distractor Hint intervention exists. Moreover, the binary transformation intervention cannot be combined with other interventions.
Please note that the instructions will change depending on the interventions. The detailed instructions are presented in 8.
Table 7: The BOAT constraints are represented in a matrix where each cell indicates whether the second intervention (column) can be applied when the first intervention (row) is introduced. A check mark (✓) denotes that the second intervention can be combined with the first one, while a black triangle ($\blacktriangle$) indicates that the second intervention is a required addition. DQ, DH, AE, OS, LR and BT represents Distractor Hint, Distractor Question, Answer Removal, Option Shuffling, Label Replacement and Binary Transformation, respectively.
First/Second Intervention DH DQ AR OS LR BT
DH $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DQ $\checkmark$ $\checkmark$ $\checkmark$
AR $\blacktriangle$ $\checkmark$ $\checkmark$
OS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LR $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BT
Table 8: The detailed instructions under different interventions. DQ, DH, AE, OS, LR and BT represents Distractor Hint, Distractor Question, Answer Removal, Option Shuffling, Label Replacement and Binary Transformation, respectively.
@@ C.2 BOAT constraints
DH If there is no correct answer in the options, please reply with N.
DQ Here are two questions and only one of them corresponds to the options. Please select the correct answer.
AR Here is a multiple choice question.
OS Here is a multiple choice question.
LR Here is a multiple choice question.
BT The following are true/false questions. If the answer is correct, please reply with T, otherwise reply with F.
Appendix D Proof of Proposition
D.1 Proof of Proposition 3.1
@@ Appendix D Proof of Proposition
Appendix E Intervened Examples